ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: Primary 47 H 10; Secondary 54 H 25 S. Sadiq Basha and P. Veeramani, Department of Mathematics Indian Institute of Technology, Madras Chennai-600036 (India) Best Proximity Pair Theorems for Multifunctions with Open Fibres Let A and B be non-empty subsets of a locally convex topological vector space with a continuous seminorm p. Given a multifunction T : AtoB with open fibres, best proximity pair theorems furnishing the sufficient conditions for the existence of an element xinA such that d(x, T x) = d(A, B) ( distance with respect to the seminorm p ) are proved.